Assume that product Alpha and product Beta are both priced at $1 per unit and that Ellie has $20 to spend on Alpha and Beta. She buys 8 units of Alpha and 12 units of Beta. The marginal utility of Alpha is 40 and the marginal utility of Beta is 20. This indicates that:
1
Expert's answer
2015-03-04T07:54:41-0500
If product Alpha and product Beta are both priced at $1 per unit and thatSalem has $20 to spend on Alpha and Beta. And he buys 8 units of Alpha and 12 units of Beta. The marginal utility of Alpha is 40 and the marginal utility of Beta is 20. This indicates that: d - in order to maximize utility, Salem should buy more of Alpha and less of Beta, because MUa/Pa = MUb/Pb when utility is maximized. In our case MUa/Pa = 40/8 = 5, MUb/Pb = 20/12 = 1.66, so MUa/Pa >MUb/Pb.